Bullseye Company manufactures dartboards. Its standard cost information follows: Standard Quantity Standard Price (Rate) Standard Unit Cost Direct materials (cork board) 2.5 sq. ft. $ 2.00 per sq. ft. $ 5.00 Direct labor 1 hrs. $ 14.00 per hr. 14.00 Variable manufacturing overhead (based on direct labor hours) 1 hrs. $ 0.50 per hr. 0.50 Fixed manufacturing overhead ($40,000 ÷ 160,000 units) 0.25 Bullseye has the following actual results for the month of September: Number of units produced and sold 140,000 Number of square feet of corkboard used 360,000 Cost of corkboard used $ 756,000 Number of labor hours worked 148,000 Direct labor cost $ 1,938,800 Variable overhead cost $ 72,000 Fixed overhead cost $ 50,000 Required: 1. Calculate the direct materials price, quantity, and total spending variances for Bullseye. (Do not round your intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable.)
Answer #1

Material price variance = (Standard price-actual price)actual quantity = (2*360000-756000) = 36000 U

Material quantity variance = (Standard quantity-actual quantity)Standard price = (140000*2.5-360000)*2 = 20000 U

Total spending variance = Standard material Cost-Actual material cost = (140000*5-756000) = 56000 U